ye 12 ut r9 selo Fig. P6-4 Fig. P6-5
P6-5. In cach case user-yeopsi and, (of the ) stirrups) 60,000 psi. 6-2, 6-3, 6-4, and 6-5 Compute for the cross sections shown in Figs. P6-2, P6-3, P6-4, and
ye 12" ut r9" selo" Fig. P6-4 Fig. P6-5
Answer #1

6a)

Width of web of concrete=5"

Effective depth=19"

Grade if concrete=4000 psi

Shear strength of concrete alone=0.75*2*sqrt(4000)*5*19/1000=9 kips

Grade of steel=60 ksi

Number of legs of #3 stirrups=2

Area of stirrups =2*0.11=0.22 in²

Spacing of stirrups=10"

Design strength of stirrups=0.75*60*0.22*19/10=28.7

Total design shear strength of section=9+28.7=37.7 kips

b)

Width of web=6+6=12 in

Effective depth=19.5 in

Shear strength of concrete=0.75*2*sqrt(4000)*12*19.5/1000=22.2 kips

Area of stirrups=2*0.2=0.4 in²

Spacing=12 in

Desugn strength of stirrups=0.75*60*0.4*19.5/12=29.25 kips

TotsT strength of section=22.2+29.25=51.45 kips